# OptiPMB: Enhancing 3D Multi-Object Tracking with Optimized Poisson Multi-Bernoulli Filtering

### v1

@@ Abstract
Index Terms:
Autonomous driving, 3D multi-object tracking, random finite set, Poisson multi-Bernoulli, Bayesian filtering.
@@ I Introduction
Accurate and reliable 3D multi-object tracking is essential for autonomous driving systems to enable robust perception, navigation, and planning in complex dynamic environments. Although deep learning has recently driven the development of many learning-based tracking methods [1, 2, 3, 4, 5, 6, 7], model-based approaches continue to attract significant attention due to their simplicity and data sample efficiency. In model-based tracking, a common design paradigm is to employ random vector (RV)-based Bayesian filtering within the tracking-by-detection (TBD) framework [8, 9, 10, 11, 12, 13, 14, 15]. Specifically, pre-trained detectors provide object bounding boxes, which serve as inputs for computing similarity scores between detections and existing objects to form an affinity matrix. Assignment algorithms, such as greedy matching and the Hungarian algorithm [16], then establish detection-to-object associations. Bayesian filters (e.g., the Kalman filter) finally update the object state vectors according to the association result. However, because RV-based Bayesian filters estimate each object’s state vector individually, these methods often depend on complex heuristic data association steps and track management schemes to effectively track multiple objects. For example, while methods in [9, 11, 12, 13, 14, 15] achieve state-of-the-art performance on large-scale autonomous driving datasets like nuScenes [17] and KITTI [18], they employ multi-stage data association and counter/score-based track management to address the inherent limitation of RV-based Bayesian filters.
@@ II-A1 TBD with RV-Based Bayesian filter
AB3DMOT [8] establishes an early open-source 3D MOT baseline on autonomous driving datasets by utilizing the TBD framework, which performs tracking using estimated object bounding boxes from pre-trained object detectors as inputs. It employs the Hungarian assignment algorithm [16] for detection-to-object association and the Kalman filter to estimate each object’s state vector, creating a simple model-based tracking paradigm that combines TBD with RV-based Bayesian filters. To further enhance tracking performance, subsequent works within this paradigm employ multi-stage data association [10, 11, 13, 14, 15], utilize different features to evaluate the affinity between objects and detections [30, 31], and combine multiple sensor modalities [12, 9, 32, 33, 34, 35]. Other data association approaches, such as joint probabilistic data association [36, 37] and multiple hypothesis tracking [38], have been applied within this paradigm as well.
@@ II-B1 TBD with Learning-Assisted Data Association
The classical strategy to enhance model-based 3D MOT with neural networks is employing learning-assisted data association. In this approach, feature embeddings captured by neural networks serve as an additional cue to construct the affinity matrix, which is subsequently processed by conventional data association methods, such as greedy matching [45, 1, 46, 3, 47] and multi-stage Hungarian algorithm [2, 48]. Object states are then estimated by the Kalman filter based on association results. Similar strategies are also applied to RFS-based filters [49, 50], where learned embeddings complement the belief propagation messages to enhance the probabilistic data association.
@@ III-A Basic Concept and Notation
In autonomous driving applications, raw observations of the objects (e.g., camera images, LiDAR and radar point clouds) are often preprocessed by detectors to obtain bird-eye’s-view (BEV) or 3D bounding boxes, enabling the trackers to perform 3D MOT using these detections under the TBD framework. Since TBD-based 3D MOT methods were proven to be simple and effective in many previous studies, we employ the TBD framework in this paper and assume that the measurement of an object is a bounding box detection. At time step $k$, assume that the objects are observed by a set of measurements $\mathbf{Z}_{k}=\{\mathbf{z}_{k}^{1},\mathbf{z}_{k}^{2},...,\mathbf{z}_{k}^{M_{k}}\}\in\mathcal{F}(\mathbb{R}^{n_{z}})$. Each measurement $\mathbf{z}_{k}\in\mathbb{R}^{n_{z}}$ is a bounding box detected from raw observations which contains measured object states such as position, velocity, size, and orientation. The collection of measurement sets from time step $1$ to $k$ is denoted by $\mathbf{Z}_{1:k}$.In practical scenarios, some objects can be misdetected, and the set of measurements $\mathbf{Z}$ often consists of not only true object detections but also false detections (clutter), as shown in Fig. 1. The ambiguous relation between the objects and the measurements increases the complexity of 3D MOT, while the RFS-based Bayesian filtering aims to provide an effective and rigorous solution to the multi-object state estimation problem.
@@ III-B Key Random Processes
A Bernoulli RFS $\mathbf{X}$ contains a single element with probability $r$ or is empty with probability $1-r$. The probability density function of $\mathbf{X}$ is given by
@@
$f^{\mathrm{mb}}(\mathbf{X})=\begin{cases}\sum_{\uplus_{i\in\mathbf{I}}\mathbf{X}^{i}=\mathbf{X}}\prod_{i\in\mathbf{I}}f^{\mathrm{ber}}(\mathbf{X}^{i})&|\mathbf{X}|\leq N\\ 0&|\mathbf{X}|>N\\ \end{cases}$ (3)
@@ III-C Bayesian Filtering with Conjugate Prior Densities
where $g(\mathbf{X}_{k}|\mathbf{X}_{k-1})$ is the multi-object state transition density. Under conventional MOT assumptions of the RFS framework [27, 28, 21], an object with state $\mathbf{x}_{k-1}$ survives from time step $k-1$ to $k$ with probability $p_{\mathrm{s}}(\mathbf{x}_{k-1})$, and its state transits with $g(\mathbf{x}_{k}|\mathbf{x}_{k-1})$. The newborn objects are modeled by a PPP RFS with intensity $\lambda^{\mathrm{b}}(\cdot)$.
@@ IV-A Single-Object State Transition and Measurement Model
$\mathbf{x}_{\mathrm{M}}$ $=[x,y,v,\phi,\omega,a]^{\mathrm{T}}$
$\mathbf{x}_{\mathrm{A}}$ $=[\alpha,\beta,\gamma,\zeta]^{\mathrm{T}}$
$\mathbf{x}_{\mathrm{U}}$ $=[c,t]^{\mathrm{T}}.$
Here, the motion state $\mathbf{x}_{\mathrm{M}}$ comprises the object’s position on the BEV plane $(x,y)$, velocity $v$, heading angle $\phi$, turn rate $\omega$, and acceleration $a$. The auxiliary state $\mathbf{x}_{\mathrm{A}}$ includes the length $\alpha$, width $\beta$, and height $\gamma$ of the object bounding box as well as $\zeta$, the object’s position along the $z$-axis. The time-invariant state $\mathbf{x}_{\mathrm{U}}$ contains the object class $c$ and track ID $t$. Finally, $x_{\mathrm{cnt}}$ records the number of consequent misdetections for this object, $x_{\mathrm{len}}$ records the number of time steps the object has survived, $x_{s}$ is a confidence score utilized in certain evaluation metrics. It is worth noting that only the motion state $\mathbf{x}_{\mathrm{M}}$ is estimated by Bayesian filtering recursions, while the non-motion states $(\mathbf{x}_{\mathrm{A}},\mathbf{x}_{\mathrm{U}},x_{\mathrm{cnt}},x_{\mathrm{len}},x_{s})$ are processed by a light-weight filter (see Section IV-F for details).
@@
$\mathbf{x}_{\mathrm{M},{k}}=\Gamma_{\mathrm{CTRA}}(\mathbf{x}_{\mathrm{M},k-1})+\mathbf{v}_{k-1}$ (7)
where $\Gamma_{\mathrm{CTRA}}(\cdot)$ denotes the nonlinear dynamic function of the CTRA model (see [65, Section III-A]), $\mathbf{v}_{k-1}\sim\mathcal{N}(\mathbf{0},Q_{k-1})$ is Gaussian-distributed process noise. Unscented transform (UT) [66] is applied to handle the nonlinear functions, and we have the motion state transition density
@@
$\mathbf{z}_{\mathrm{M}}$ $=[z_{x},z_{y},z_{vx},z_{vy},z_{\phi}]^{\mathrm{T}}$
$\mathbf{z}_{\mathrm{A}}$ $=[z_{\alpha},z_{\beta},z_{\gamma},z_{\zeta}]^{\mathrm{T}}.$
Here, $\mathbf{z}_{\mathrm{M}}$ is the observation of the object’s motion state, while $\mathbf{z}_{\mathrm{A}}$ denotes the observation of the object’s auxiliary state. The motion measurement model is given by
$\mathbf{z}_{\mathrm{M},k}=\mathrm{H}(\mathbf{x}_{\mathrm{M},k})+\mathbf{w}_{k}=\begin{bmatrix}x_{k}\\ y_{k}\\ v_{k}\cos\phi_{k}\\ v_{k}\sin\phi_{k}\\ \phi_{k}\end{bmatrix}+\mathbf{w}_{k}$ (10)
@@ IV-B Basic Framework: Bayesian Filtering with PMB Density
where the multi-object state is modeled as a union of two disjoint subsets, i.e., $\mathbf{X}^{\mathrm{u}}_{k}\uplus\mathbf{X}^{\mathrm{d}}_{k}=\mathbf{X}_{k}$. The PPP RFS $\mathbf{X}^{\mathrm{u}}_{k}$ denotes the state of potential objects which have never been detected, while the MB RFS $\mathbf{X}^{\mathrm{d}}_{k}$ denotes the state of potential objects which have been detected at least once. This partition of detected and undetected objects enables an efficient hypotheses management of potential objects [25, 26, 27].
According to the definitions in Section III-B, the posterior PMB density at time step $k-1$, $f(\mathbf{X}_{k-1}|\mathbf{Z}_{1:k-1})$, can be fully determined by a set of parameters
@@
  • First-Time Detection Hypotheses:$\{l^{1,3},l^{2,4},l^{3,5}\}$, indicating that the measurements originate from newly detected objects or clutter, corresponding to new potential objects $\{\mathbf{x}^{3},\mathbf{x}^{4},\mathbf{x}^{5}\}$.
@@
As illustrated in Fig. 2, at the subsequent time step $k$, the PMB density $f(\mathbf{X}_{k-1}|\mathbf{Z}_{1:k-1})$ is first propagated through the prediction step and then updated using the preprocessed object detections. The situation that previously undetected potential objects remain undetected at the current time step is also considered within this pipeline (see Section 28) alongside the three types of local hypotheses. Following the PMB prediction and update, the data association module determines the optimal global data association hypothesis, which is then utilized as protocol for the pruning module to eliminate redundant PPP and Bernoulli components from the PMB density. Next, the non-motion states (including track IDs) of detected potential objects are determined by a light-weight filter. Finally, the estimated posterior $f(\mathbf{X}_{k}|\mathbf{Z}_{1:k})$ is processed by the track extraction module to generate the object tracks. The details of our OptiPMB tracker are explained in the following subsections, where the time subscript $k|k-1$ is abbreviated as $+$ for simplicity. Pseudo code for the complete OptiPMB recursion is provided in Algorithm 1.
@@
4Note: The subscript $+$ is the abbreviation of $k|k-1$. All detection probability $p_{\mathrm{d}}(\mathbf{x}_{+})$ are adaptively determined (see Section IV-G1).
@@ 56                   Compute $r_{k}^{m,n},p_{k}^{m,n},w_{k}^{m,n}$ using (37);
Algorithm 1 Pseudo Code of OptiPMB (Part I)
@@ IV-D1 MB Update for Misdetection and Detection Hypotheses
$\begin{cases}r=\frac{r_{+}^{n}(1-p_{\mathrm{d}}(\mathbf{x}_{+}^{n}))}{1-r_{+}^{n}+r_{+}^{n}(1-p_{\mathrm{d}}(\mathbf{x}_{+}^{n}))}\\ p(\mathbf{x})=p_{+}^{n}(\mathbf{x})\\ \end{cases}$ (17)
@@
1
36 $\hat{N}_{k}^{\mathrm{d}}\leftarrow N_{k-1}^{\mathrm{d}}+|\mathbf{Z}_{k}|$, $\hat{\mathbf{I}}_{k}\leftarrow\uplus_{n=1}^{\hat{N}_{k}^{\mathrm{d}}}\hat{\mathbf{L}}_{k}^{n}$;
39
40/* Data association and PMB pruning (see Section IV-E and Section IV-H1) */
41 From $\{w_{k}^{m,n}\}$, construct the cost matrix $W$;
44 /* Light-weight filtering (see Section IV-F) */
45 Update non-motion states with the light-weight filter;
47
48/* Track extraction (see Section IV-H2) */
Algorithm 2 Pseudo Code of OptiPMB (Part II)
@@ IV-D2 MB Update for First-Time Detection Hypotheses
$\begin{cases}w=-\ln[e+\lambda^{\mathrm{c}}(\mathbf{z}_{k}^{m})]\\ r=e/[e+\lambda^{\mathrm{c}}(\mathbf{z}_{k}^{m})]\\ p(\mathbf{x})=\frac{1}{e}\sum_{j\in\mathbf{J}^{m}_{k}}e^{j}p^{j}(\mathbf{x})\\ e=\sum_{j\in\mathbf{J}^{m}_{k}}e^{j}\end{cases}$ (23)
@@
$\begin{cases}e^{j}=\mu_{k-1}^{\mathrm{u},j}p_{\mathrm{d}}(\mathbf{x}_{+}^{\mathrm{u},j})\mathcal{N}(\mathbf{z}_{xy,k}^{m};\hat{\mathbf{z}}^{\mathrm{u},j}_{k},S_{k}^{\mathrm{u},j})\\ p^{j}(\mathbf{x})=\mathcal{N}(\mathbf{x};\mathbf{m}_{k}^{m,j},P_{k}^{m,j})\end{cases}$ (24)

### v2

@@ I Introduction
Accurate and reliable 3D multi-object tracking is essential for autonomous driving systems to enable robust perception, navigation, and planning in complex dynamic environments. Although deep learning has recently driven the development of many learning-based tracking methods [1, 2, 3, 4, 5, 6, 7], model-based approaches continue to attract significant attention due to their simplicity and data sample efficiency. In model-based tracking, a common design paradigm is to employ random vector (RV)-based Bayesian filtering within the tracking-by-detection (TBD) framework [8, 9, 10, 11, 12, 13, 14, 15, 16]. Specifically, pre-trained detectors provide object bounding boxes, which serve as inputs for computing similarity scores between detections and existing objects to form an affinity matrix. Assignment algorithms, such as greedy matching and the Hungarian algorithm [17], then establish detection-to-object associations. Bayesian filters (e.g., the Kalman filter) finally update the object state vectors according to the association result. However, because RV-based Bayesian filters estimate each object’s state vector individually, these methods often depend on complex heuristic data association steps and track management schemes to effectively track multiple objects. For example, while methods in [9, 11, 12, 13, 14, 15, 16] achieve state-of-the-art performance on large-scale autonomous driving datasets like nuScenes [18] and KITTI [19], they employ multi-stage data association and counter/score-based track management to address the inherent limitation of RV-based Bayesian filters.
@@ II-A1 TBD with RV-Based Bayesian filter
AB3DMOT [8] establishes an early open-source baseline for 3D MOT on autonomous driving datasets by adopting the TBD framework, which performs tracking with object bounding boxes estimated by pre-trained object detectors. It employs the Hungarian algorithm [17] for detection-to-object association and the Kalman filter (KF) to estimate object state vectors, thereby forming a simple model-based paradigm that integrates TBD with RV-based Bayesian filters. Building upon this paradigm, subsequent works have enhanced tracking performance by introducing multi-stage data association [10, 11, 13, 14, 15, 16], exploiting diverse features to measure affinities between detections and objects [25, 26], and incorporating multiple sensing modalities [12, 9, 27, 28, 29, 30, 16]. For example, MCTrack [15] performs two-stage matching on both the bird-eye’s-view (BEV) plane and the range view plane, thereby reducing false positives and ID switch errors when using low-quality detections. EMMS-MOT [16] leverages spatial synergies between 3D and 2D detection boxes by jointly estimating 3D and 2D motion states with KF and adopting a three-stage data association strategy. Moreover, alternative association methods, including joint probabilistic data association [31, 32] and multiple hypothesis tracking [33], have also been explored within this paradigm.
@@ II-B1 TBD with Learning-Assisted Data Association
The classical strategy to enhance model-based 3D MOT with neural networks is employing learning-assisted data association. In this approach, feature embeddings captured by neural networks serve as an additional cue to construct the affinity matrix, which is subsequently processed by conventional data association methods, such as greedy matching [39, 1, 40, 3, 41] and multi-stage Hungarian algorithm [2, 42, 43]. Object states are then estimated by the Kalman filter based on association results. For example, DINO-MOT [43] leverages the DINOv2 encoder to extract visual features of pedestrians and refines the association results by computing cosine similarities between embeddings. Similar strategies are also applied to RFS-based filters [44, 45], where learned embeddings complement the belief propagation messages to enhance the probabilistic data association.
@@ III-A Basic Concept and Notation
In autonomous driving applications, raw observations of the objects (e.g., camera images, LiDAR and radar point clouds) are often preprocessed by detectors to obtain bounding boxes, enabling the trackers to perform MOT using these detections under the TBD framework. Since TBD-based 3D MOT methods were proven to be simple and effective in many previous studies, we employ the TBD framework in this paper and assume that the measurement of an object is a bounding box detection. At time step $k$, assume that the objects are observed by a set of measurements $\mathbf{Z}_{k}=\{\mathbf{z}_{k}^{1},\mathbf{z}_{k}^{2},...,\mathbf{z}_{k}^{M_{k}}\}\in\mathcal{F}(\mathbb{R}^{n_{z}})$. Each measurement $\mathbf{z}_{k}\in\mathbb{R}^{n_{z}}$ is a bounding box detected from raw observations which contains measured object states such as position, velocity, size, and orientation. The collection of measurement sets from time step $1$ to $k$ is denoted by $\mathbf{Z}_{1:k}$.In practical scenarios, some objects can be misdetected, and the set of measurements $\mathbf{Z}$ often consists of not only true object detections but also false detections (clutter), as shown in Fig. 1. The ambiguous relation between the objects and the measurements increases the complexity of 3D MOT, while the RFS-based Bayesian filtering aims to provide an effective and rigorous solution to the multi-object state estimation problem.
@@ III-B Key Random Processes
A Bernoulli RFS $\mathbf{X}$ contains a single element with probability $r$ or is empty with probability 1r1-r. The probability density function of $\mathbf{X}$ is given by
@@
[…]
@@ III-C Bayesian Filtering with Conjugate Prior Densities
where $g(\mathbf{X}_{k}|\mathbf{X}_{k-1})$ is the multi-object state transition density. Under conventional MOT assumptions of the RFS framework [60, 62, 22, 61], an object with state $\mathbf{x}_{k-1}$ survives from time step k1k-1 to $k$ with probability $p_{\mathrm{s}}(\mathbf{x}_{k-1})$, and its state transits with $g(\mathbf{x}_{k}|\mathbf{x}_{k-1})$. The newborn objects are modeled by a PPP RFS with intensity $\lambda^{\mathrm{b}}(\cdot)$.
@@ IV-A Single-Object State Transition and Measurement Model
where $\mathbf{x}_{\mathrm{M}}=[x,y,v,\phi,\omega,a]^{\mathrm{T}}$, $\mathbf{x}_{\mathrm{A}}=[\alpha,\beta,\gamma,\zeta]^{\mathrm{T}}$, and $\mathbf{x}_{\mathrm{U}}=[c,t]^{\mathrm{T}}$. The motion state $\mathbf{x}_{\mathrm{M}}$ comprises the object’s position on the BEV plane $(x,y)$, velocity $v$, heading angle $\phi$, turn rate $\omega$, and acceleration $a$. The auxiliary state $\mathbf{x}_{\mathrm{A}}$ includes the length $\alpha$, width $\beta$, and height $\gamma$ of the object bounding box as well as $\zeta$, the object’s position along the $z$-axis. The time-invariant state $\mathbf{x}_{\mathrm{U}}$ contains the object class $c$ and track ID $t$. Finally, $x_{\mathrm{cnt}}$ records the number of consequent misdetections for this object, $x_{\mathrm{len}}$ records the number of time steps the object has survived, $x_{s}$ is a confidence score utilized in certain evaluation metrics. It is worth noting that only the motion state $\mathbf{x}_{\mathrm{M}}$ is estimated by Bayesian filtering recursions, while the non-motion states $(\mathbf{x}_{\mathrm{A}},\mathbf{x}_{\mathrm{U}},x_{\mathrm{cnt}},x_{\mathrm{len}},x_{s})$ are processed by a light-weight filter (see Section IV-F for details).
@@
$\mathbf{x}_{\mathrm{M},{k}}=\Gamma_{\mathrm{CTRA}}(\mathbf{x}_{\mathrm{M},k-1})+\mathbf{v}_{k}$ (7)
where $\Gamma_{\mathrm{CTRA}}(\cdot)$ denotes the nonlinear dynamic function of the CTRA model (see [63, Section III-A]), $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},Q_{k})$ is Gaussian-distributed process noise. Unscented transform (UT) [64] is applied to handle the nonlinear functions, and we have the motion state transition density
@@
where $\mathbf{z}_{\mathrm{M}}=[z_{x},z_{y},z_{vx},z_{vy},z_{\phi}]^{\mathrm{T}}$ and $\mathbf{z}_{\mathrm{A}}=[z_{\alpha},z_{\beta},z_{\gamma},z_{\zeta}]^{\mathrm{T}}$. Here, $\mathbf{z}_{\mathrm{M}}$ is the observation of the object’s motion state, while $\mathbf{z}_{\mathrm{A}}$ denotes the observation of the object’s auxiliary state. The motion measurement model is given by
$\mathbf{z}_{\mathrm{M},k}$ $=\mathrm{H}(\mathbf{x}_{\mathrm{M},k})+\mathbf{w}_{k}$ (10)
$=[x_{k},y_{k},v_{k}\cos\phi_{k},v_{k}\sin\phi_{k},\phi_{k}]^{\mathrm{T}}+\mathbf{w}_{k}$
@@ IV-B Basic Framework: Bayesian Filtering with PMB Density
where the multi-object state is modeled as a union of two disjoint subsets, i.e., $\mathbf{X}^{\mathrm{u}}_{k}\uplus\mathbf{X}^{\mathrm{d}}_{k}=\mathbf{X}_{k}$. The PPP RFS $\mathbf{X}^{\mathrm{u}}_{k}$ denotes the state of potential objects which have never been detected, while the MB RFS $\mathbf{X}^{\mathrm{d}}_{k}$ denotes the state of potential objects which have been detected at least once. This partition of detected and undetected objects enables an efficient hypotheses management of potential objects [61, 60].
According to the definitions in Section III-B, the posterior PMB density at time step k1k-1, $f(\mathbf{X}_{k-1}|\mathbf{Z}_{1:k-1})$, can be fully determined by a set of parameters
@@
[…]
@@
As illustrated in Fig. 2, at the subsequent time step $k$, the PMB density $f(\mathbf{X}_{k-1}|\mathbf{Z}_{1:k-1})$ is first propagated through the prediction step and then updated using the preprocessed object detections. The situation that previously undetected potential objects remain undetected at the current time step is also considered within this pipeline (see Section 28) alongside the three types of local hypotheses. Following the PMB prediction and update, the data association module determines the optimal global data association hypothesis, which is then utilized as protocol for the pruning module to eliminate redundant PPP and Bernoulli components from the PMB density. Next, the non-motion states (including track IDs) of detected potential objects are determined by a light-weight filter. Finally, the estimated posterior $f(\mathbf{X}_{k}|\mathbf{Z}_{1:k})$ is processed by the track extraction module to generate the object tracks. The details of our OptiPMB tracker are explained in the following subsections, where the time subscript k|k1k|k-1 is abbreviated as ++ for simplicity. Pseudo code for the complete OptiPMB recursion is provided in Algorithm 1.
@@
4Note: The subscript ++ is the abbreviation of k|k1k|k-1. All detection probability $p_{\mathrm{d}}(\mathbf{x}_{+})$ are adaptively determined (see Section IV-G1).
@@ 56        Compute $r_{k}^{m,n},p_{k}^{m,n},w_{k}^{m,n}$ using (37);
60
[…]
64
65/* Data association and PMB pruning (see Section IV-E and Section IV-H1) */
66 From $\{w_{k}^{m,n}\}$, construct the cost matrix $W$;
69 /* Light-weight filtering (see Section IV-F) */
70 Update non-motion states with the light-weight filter;
72
73/* Track extraction (see Section IV-H2) */
Algorithm 1 Pseudo Code of OptiPMB
@@ IV-D1 MB Update for Misdetection and Detection Hypotheses
[…]
@@ IV-D2 MB Update for First-Time Detection Hypotheses
$\begin{cases}w=-\ln[e+\lambda^{\mathrm{c}}(\mathbf{z}_{k}^{m})]\\ r=e/[e+\lambda^{\mathrm{c}}(\mathbf{z}_{k}^{m})]\\ p(\mathbf{x})=\frac{1}{e}\sum_{j\in\mathbf{J}^{m}_{k}}e^{j}p^{j}(\mathbf{x})\end{cases}$ (23)
@@
$\begin{cases}e=\sum_{j\in\mathbf{J}^{m}_{k}}e^{j}\\ e^{j}=\mu_{+}^{\mathrm{u},j}p_{\mathrm{d}}(\mathbf{x}_{+}^{\mathrm{u},j})\mathcal{N}(\mathbf{z}_{xy,k}^{m};\hat{\mathbf{z}}^{\mathrm{u},j}_{k},S_{k}^{\mathrm{u},j})\\ p^{j}(\mathbf{x})=\mathcal{N}(\mathbf{x};\mathbf{m}_{k}^{m,j},P_{k}^{m,j})\end{cases}$ (24)